Determine the discounted value of an investment with maturity value of $12,000, if it is discounted 5 years before maturity at 8% compounded quarterly.

Select one:

a. $8590.48

b. $8075.66

c. $8384.39

d. $8129.37

e. $8152.69
Answer #1

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

12000=P(1+0.08/4)^(4*5)

P=12000/(1+0.08/4)^(4*5)

=$12000*0.672971333

which is equal to

=$8075.66(Approx).